OpenStudy (anonymous):

May has 10 fewer marbles than Albert May has 5 more marbles than Jenny Albert gives 4 marbles to Jenny May gives antoher 6 marbles to Jenny Jenny has 13 marbles now how much may and ablert have at first

OpenStudy (primeralph):

M = A-10 M = 5+J A-->A-4, M-->M-6; while J-->J+4+6 = 13 Originally, J = 13-10 = 3. M = 5+3 = 8. A = 10+M = 18.

OpenStudy (anonymous):

This is exactly the answer we got and my son's teacher said it incorrect.

OpenStudy (primeralph):

That's the answer to the problem; unless some information is missing.